OpenStudy (anonymous):

Could someone please help me with this problem? I am unfamiliar with point-slope/slope intercept. So I am unsure how to go about the conversion. Thank you!! Rewrite the equation of a line in point-slope form to slope intercept form: y-2 = 2/3 (6x+9)

OpenStudy (jdoe0001):

hmm lemme redo that \(\bf y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad\\ \qquad \uparrow\\ \textit{point-slope form} \\ \quad \\ y={\color{green}{ m}}x+y-intercept\\ \qquad \uparrow \\ \textit{slope-intercept form}\) so in short notice the number to the right-side of the "y", that's the y-coordinate notice the number to the right-side of the "x", that's the x-coordinate "m" is the slope, or the number in front of the parentheses so you expand all those and simplify them to make it look like the slope-intercept form so you'd end up with y = "some number" x + "some other number" like for example y = (3/5) x + 11 or so

OpenStudy (jdoe0001):

you could just start it up by simply solving the equation for "y", so the "y' would be left on the left-hand-side by itself

OpenStudy (jdoe0001):

lemme make a quick example say the point-slope form is y - 3 = 5(x+1) y-3 = 5x+5 y = 5x+5 +3 y = 5x+8 <---- slope-intercept form